OpenStudy (anonymous):

simplify -32/85 / 50/51= -32/85 * 50/51

OpenStudy (ganpat):

Hint: ( a/b ) / (c/d) = (a/b) * (d/c)

OpenStudy (anonymous):

-32/85 / 50/51 -32/85 * 50/51 -32/17 * 10/51 -320/17*51 -320/867 -32/85 / 50/51= -320/867

OpenStudy (ganpat):

good, now use the formula, what i gave u on LHS..

OpenStudy (anonymous):

huh

OpenStudy (ganpat):

what exactly is the question , can u write again, plz ?

OpenStudy (ganpat):

@amber517 : SOmething is wrong.. -32/85 / 50/51= -32/85 * 51/50..

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

what is wrong

OpenStudy (ganpat):

look, i believe the given question is -32/85 / 50/51.. right ??

OpenStudy (anonymous):

yes

OpenStudy (ganpat):

so its conversion is -32/85 * 51/50.. You have written wrong.. check the RHS of posted question ... you have written 50/51, but it should be 51/50..

OpenStudy (anonymous):

ooo ok

OpenStudy (anonymous):

can this be simplifyed